# The Alexander polynomial of twisted torus knots

### v1

@@ 4. Degree of the Alexander polynomial of Twisted Torus Knots
$B=t^{pq+(r-1)rs}-t^{Q^{\prime}q+(r-1)rs}$
We obtain $\tilde{X}Y-X\tilde{Y}=\tilde{X}B-\tilde{Y}A$. Since $|rs|>Q_{i}q$ and $|rs|>\overline{k_{i}}p$ for all $i$, then the highest powers in $\tilde{X}B$ and $\tilde{Y}A$ are $pq+(r-1)rs+k_{1}p$ and $pq+(r-1)rs+Q_{1}q$, respectively. Similarly the lowest powers in $\tilde{X}B$ and $\tilde{Y}A$ are $Q^{\prime}q+(r-1)rs+\overline{k_{r-1}}p+(r-2)rs+rs$ and $k^{\prime}p+(r-1)rs+Q_{r-1}q+(r-2)rs+rs$, respectively. Since $k_{1}p<Q_{1}q$, and by Lemma 9, the highest and the lowest powers if $t$ in $\tilde{X}Y-X\tilde{Y}$ are $pq+(r-1)rs+Q_{1}q$ and $k^{\prime}p+(r-1)rs+Q_{r-1}q+(r-2)rs+rs$, respectively. Hence,
$\text{deg}(\Delta_{T(p,q;r,s)})=(p-1)(q-1)+(Q_{1}-Q^{\prime}-Q_{r-1})q-(r-1)rs.$
Now, consider the case when $10<m<r-$ 1. Since $|rs|>Q_{i}q$, $|rs|>\overline{k_{i}}p$ for all $i$, the highest and lowest powers of $t$ in $\tilde{X}B$ and $A\tilde{Y}$ of Equation (7) of Proposition 11 are $Q_{m+1}q+mrs+k_{1}p$, $Q_{r-1}q+(r-2)rs+\overline{k_{m}}p+(m-1)rs+2rs$ and $\overline{k_{m+1}}p+mrs+Q_{1}q$, $\overline{k_{r-1}}p+(r-2)rs+Q_{m}q+(m-1)rs+2rs$, respectively. Since $Q_{i}q>\overline{k_{i}}p$ for all $i$, the highest and lowest power of $t$ in $\tilde{X}Y-X\tilde{Y}$ are $\overline{k_{m+1}}p+mrs+Q_{1}q$ and $\overline{k_{r-1}}p+(r-2)rs+Q_{m}q+(m-1)rs+2rs$, respectively. Therefore,
@@
Figure 6. An isotopy of $T(9,8;3,-1)$ to a diagram that produce a Seifert surface with the minimal genus via Seifert algorithm.
The degree of twisted torus knot $T(p,p-1;3,-1)$ provides a sharp lower bound for the genus of the knot. By applying a similar isotopy as shown in Figure 6 and using the Seifert algorithm on the resulting diagram, we obtain a Seifert surface of $T(p,p-1;3,-1)$ with a genus that coincides with the half of the degree of the Alexander polynomial of the knot. Hence, the genus of the knot is equal to the genus of the surface. Since the final diagram of the knot is, in fact, the closure of a positive braid, one can also show that the Seifert surface obtained from this diagram yields the genus of the knot.
@@ 5. Twisted Torus Knots Not Admitting Lens Space Surgery
Although our formula for the Alexander polynomial of a twisted torus knot involves integers related to modular arithmetic, it can be simplified by focusing specific families of twisted torus knots. We present an example here to illustrate this simplification and to prove Theorem 3 concerning lens space surgery.
Consider a twisted torus knot of type $T(2n+1,n;r,s)$. In this case, note that $[n^{-1}]=2n-1$. Therefore,
$Q=\{\left[i(2n-1)\right]|i=0,1,\dots,r-1\}=\{0,2n-1,2n-3,\dots,2n-(2r-3)\},$
and so
$Q_{0}=0\quad\text{and}\quad Q_{i}=2(n-r)+1+2i\quad\text{for}\quad i=1,\dots,r-1.$
We also have
$R=\{[-i(2n-1)]|i=1,2,\dots,n\}=\{2,4,\dots,2n\},$
and hence
$k_{1}=n-r+1\quad\text{and}\quad k_{2}=k_{3}=\dots=k_{r}=1$
with
$\overline{k_{i}}=n-r+i\quad\text{for}\quad i=1,\dots r.$
Since $Q^{\prime}=[r(2n-1)]=2(n-r)+1$, we obtain
$m=0\quad\text{and}\quad\overline{k^{\prime}}=n-r.$
Therefore, we have
$X(t)=$ $1-(1-t^{rs})\sum_{i=1}^{r-1}t^{(n-r+i)(2n+1)+(i-1)rs}-t^{(2n+1)n+(r-1)rs},$
$\tilde{X}(t)=$ $1-t^{(2n+1)(n-r)},$
$Y(t)=$ $1-(1-t^{rs})\sum_{i=1}^{r-1}t^{\left(2(n-r)+1+2i\right)n+(i-1)rs}-t^{(2n+1)n+(r-1)rs},\quad\text{and}$
$\tilde{Y}(t)=$ $1-t^{(2(n-r)+1)n}.$
The Alexander polynomial of $T(2n+1,n;r,s)$ as follows:
$\Delta_{T(2n+1,n;r,s)}(t)=\frac{(1-t)(\tilde{X}(t)Y(t)-X(t)\tilde{Y}(t))}{(1-t^{2n+1})(1-t^{n})(1-t^{r})}$
Now, assume $n-2\geq r\geq 2$, and $s\geq 2$. We claim that the coefficient of $t^{2n^{2}-2nr+3n+1}$, which is the eighth lowest power term of $\Delta_{T(2n+1,n;r,s)}(t)$, is 2. To determine this coefficient of $t^{2n^{2}-2nr+3n+1}$, we may disregard all terms in $\tilde{X}(t)Y(t)-X(t)\tilde{Y}(t)$ whose power is higher than $12n^{2}-2nr+3n+$ 1. Thus, under the assumptions for $n$, $r$, and $s$, $\tilde{X}(t)Y(t)-X(t)\tilde{Y}(t)$ only includes the following terms with powers less than or equal to this value:
$-t^{(2n+1)(n-r)}+t^{(2(n-r)+1)n}+t^{(n-r+1)(2n+1)}-t^{(2(n-r)+3)n}$
@@
$(\tilde{X}(t)Y(t)-X(t)\tilde{Y}(t))(1-t)(1+t^{(2n+1)}+t^{2(2n+1)}+\dots)(1+t^{n}+t^{2n}+\dots)(1+t^{r}+t^{2r}+\dots)$
of the Alexander polynomial, all terms with the power $12n^{2}-2nr+3n+$ 1 are as follows:
$-t^{(2n+1)(n-r)+(2n+1)+r},\quad t^{(2n+1)(n-r)+1+2n+r},\quad-t^{(2(n-r)+1)n+1+2n},\quad t^{(2(n-r)+1)n+(2n+1)},$
$t^{(n-r+1)(2n+1)+r},\quad t^{(2(n-r)+3)n+1}.$
Therefore, the coefficient of $t^{2n^{2}-2nr+3n+1}$ is 2, and Theorem 3 follows.

### v3

@@ 4. Degree of the Alexander polynomial of Twisted Torus Knots
$\tilde{Y}=1-(1-t^{rs})\sum_{i=1}^{r-1}t^{Q_{i}q+(i-1)rs}-t^{Q^{\prime}q+(r-1)rs}.$
We write $X=\tilde{X}-A$ and $Y=\tilde{Y}-B$, where
$A=t^{pq+(r-1)rs}-t^{\overline{k^{\prime}}p+(r-1)rs}\quad\text{and}\quad B=t^{pq+(r-1)rs}-t^{Q^{\prime}q+(r-1)rs}.$
Then $\tilde{X}Y-X\tilde{Y}=\tilde{X}B-\tilde{Y}A$. The highest powers in $\tilde{X}B$ and $\tilde{Y}A$ are $pq+k_{1}p+(r-1)rs$ and $pq+Q_{1}q+(r-1)rs$, respectively. Similarly, the lowest powers in $\tilde{X}B$ and $\tilde{Y}A$ are $Q^{\prime}q+\overline{k_{r-1}}p+2(r-1)rs$ and $Q_{r-1}q+\overline{k^{\prime}}p+2(r-1)rs$, respectively. By Lemma 9, the highest and the lowest powers in $\tilde{X}Y-X\tilde{Y}$ are $pq+Q_{1}q+(r-1)rs$ and $\overline{k^{\prime}}p+Q_{r-1}q+2(r-1)rs$, respectively. Hence,
$\text{deg}\left(\Delta_{T(p,q;r,s)}\right)=(p-1)(q-1)+(Q_{1}-Q^{\prime}-Q_{r-1})q-(r-1)rs.$
Now, consider the case when $10<m<r-$ 1. The highest powers of $t$ in $\tilde{X}B$ and $\tilde{Y}A$ of Equation (7) are $Q_{m+1}q+k_{1}p+mrs$ and $Q_{1}q+\overline{k_{m+1}}p+mrs$ respectively, and the lowest powers in $\tilde{X}B$ and $\tilde{Y}A$ are $Q_{r-1}q+\overline{k_{m}}p+(r+m-1)rs$ and $Q_{m}q+\overline{k_{r-1}}p+(r+m-1)rs$ respectively.
We claim that $Q_{1}q-k_{1}p>Q_{m+1}q-\overline{k_{m+1}}p$. Let $n_{1}:=Q_{1}q-k_{1}p$ and $n_{m+1}:=Q_{m+1}q-\overline{k_{m+1}}p$. Then, by Lemma 9, we have $Q_{1}=[n_{1}q^{-1}]$ and $Q_{m+1}=[n_{m+1}q^{-1}]$. Consider the following sequence of residues of multiples of $q^{-1}$ modulo $p$:
$a_{1}=[q^{-1}],a_{2}=[2q^{-1}],\dots,a_{r}=[rq^{-1}].$
Note that $[rq^{-1}]$ cannot be either the minimum or the maximum value in this sequence, as that would imply $m=0\text{ or }r-1$, by the definition of $m$, contradicting our assumption. Hence, $n_{1}$ corresponds to the index of the minimal value, and $n_{m+1}$ corresponds to the index of the next greater value than $[rq^{-1}]$. Moreover, since $m\neq r-1$, we must have $n_{1}\neq 1$. Examining the sequence of residues modulo $p$, we observe that the minimal value appears later in the sequence than the next greater value following $[rq^{-1}]$. This implies that $n_{1}>n_{m+1}$, proving the claim.
Similarly, we can show that $Q_{m}q-\overline{k_{m}}p<Q_{r-1}q-\overline{k_{r-1}}p$. Then, it follows from these inequalities and Equation (7) that the highest and lowest powers in $\tilde{X}Y-X\tilde{Y}$ are $Q_{1}q+\overline{k_{m+1}}p+mrs$ and $Q_{m}q+\overline{k_{r-1}}p+(r+m-1)rs$, respectively. Therefore, we obtain
@@
Figure 6. An isotopy of $T(9,8;3,-1)$ to a diagram that yields a Seifert surface of minimal genus via Seifert algorithm.
The degree of twisted torus knot $T(p,p-1;3,-1)$ provides a sharp lower bound for the genus of the knot. By applying a similar isotopy as shown in Figure 6 and using the Seifert algorithm on the resulting diagram, we obtain a Seifert surface of $T(p,p-1;3,-1)$ with a genus that coincides with half of the degree of the Alexander polynomial of the knot. Hence, the genus of the knot is equal to the genus of the surface. Since the final diagram of the knot is, in fact, the closure of a positive braid, one can also show that the Seifert surface obtained from this diagram yields the genus of the knot.
@@ 5. Twisted Torus Knots Not Admitting Lens Space Surgery
Although our formula for the Alexander polynomial of a twisted torus knot involves integers related to modular arithmetic, it can be simplified by focusing on specific families of twisted torus knots. We present an example here to illustrate this simplification and to prove Theorem 3 concerning lens space surgery.
Consider a twisted torus knot of the type $T(9+2n,7+2n;3,s)$. Recalling the computation from the example preceding Theorem 7, we have
$X(t)=$ $1-(1-t^{3s})\sum_{i=1}^{2}t^{i(3+n)(9+2n)+(i-1)3s}-t^{(9+2n)(7+2n)+6s},$
$\tilde{X}(t)=$ $1-t^{(2+n)(9+2n)},$
$Y(t)=$ $1-(1-t^{3s})\sum_{i=1}^{2}t^{i(4+n)(7+2n)+(i-1)3s}-t^{(9+2n)(7+2n)+6s},\quad\text{and}$
$\tilde{Y}(t)=$ $1-t^{(3+n)(7+2n)}.$
The Alexander polynomial of $T(9+2n,7+2n;3,s)$ as follows:
$\Delta_{T(9+2n,7+2n;3,s)}(t)=\frac{(1-t)(\tilde{X}(t)Y(t)-X(t)\tilde{Y}(t))}{(1-t^{9+2n})(1-t^{7+2n})(1-t^{3})}$
Now, assume $s\geq 2$. We claim that the coefficient of $t^{31+15n+2n^{2}}$, which is the eighth lowest power term of $\Delta_{T(9+2n,7+2n;3,s)}(t)$, is $-2$. To determine this coefficient of $t^{31+15n+2n^{2}}$, we may disregard all terms in $\tilde{X}(t)Y(t)-X(t)\tilde{Y}(t)$ whose power is higher than $31+15n+2n^{2}$. Thus, under the assumptions for $n$ and $s$, $\tilde{X}(t)Y(t)-X(t)\tilde{Y}(t)$ only includes the following terms with powers less than or equal to this value:
$-t^{18+13n+2n^{2}}+t^{21+13n+2n^{2}}+t^{(3+n)(9+2n)}-t^{(4+n)(7+2n)}$
@@
$(\tilde{X}(t)Y(t)-X(t)\tilde{Y}(t))(1-t)(1+t^{(9+2n)}+t^{2(9+2n)}+\dots)(1+t^{7+2n}+t^{2(7+2n)}+\dots)(1+t^{3}+t^{6}+\dots)$
of the Alexander polynomial, all terms with the power $31+15n+2n^{2}$ are as follows:
$t^{(18+13n+2n^{2})+1+(9+2n)+3},\quad-t^{(18+13n+2n^{2})+(7+2n)+6},\quad t^{(21+13n+2n^{2})+(7+2n)+3},$
$-t^{(21+13n+2n^{2})+1+(9+2n)},\quad-t^{(3+n)(9+2n)+1+3},\quad-t^{(4+n)(7+2n)+3}.$
Therefore, the coefficient of $t^{31+15n+2n^{2}}$ is $-2$, and Theorem 3 follows.